Yesterday evening Nic and me were in the lab. The Mode Cleaner was unlocked, but after many attempt we could fix it and we did many scans of the Y arm cavity.
Today I was not able to keep the MC locked. Koji helped me remotely, and eventually the MC locked back, but after half an hour of measurements I had to stop.
I made some more scan of the Y arm though. I also tried to do the same for the X arm, but the MC unlocked before the measurement was finished. I'll try to come back in the night.